# Knowledge Boundary of Large Language Models: A Survey

### v1

@@ Abstract
Although large language models (LLMs) store vast amount of knowledge in their parameters, they still have limitations in the memorization and utilization of certain knowledge, leading to undesired behaviors such as generating untruthful and inaccurate responses. This highlights the critical need to understand the knowledge boundary of LLMs, a concept that remains inadequately defined in existing research. In this survey, we propose a comprehensive definition of the LLM knowledge boundary and introduce a formalized taxonomy categorizing knowledge into four distinct types. Using this foundation, we systematically review the field through three key lenses: the motivation for studying LLM knowledge boundaries, methods for identifying these boundaries, and strategies for mitigating the challenges they present. Finally, we discuss open challenges and potential research directions in this area. We aim for this survey to offer the community a comprehensive overview, facilitate access to key issues, and inspire further advancements in LLM knowledge research. Corresponding author.
@@
Moxin Li1∗    Yong Zhao1∗   Yang Deng2† Wenxuan Zhang   Shuaiyi Li3   Wenya Xie4   See-Kiong Ng1   Tat-Seng Chua1 1National University of Singapore  2Singapore Management University 3The Chinese University of Hong Kong  4The Chinese University of Hong Kong, Shenzhen {limoxin,yzhao}@u.nus.eduydeng@smu.edu.sg
@@ 1 Introduction
We identify the major limitations in existing definitions of the LLM knowledge boundary. Firstly, the Know-Unknow Quadrant Yin et al. (2023); Amayuelas et al. (2024) categorizes knowledge based on the LLM’s possession and the LLM’s awareness of such knowledge, but this definition is conceptual and lacks formalization. Besides, Yin et al. (2024) introduce a formalized definition separating the influence of the prompt from the LLM’s mastery of the knowledge, yet they merely focus on the knowledge boundary of a specific LLM which lacks comprehensiveness. Additionally, some recent surveys Li et al. (2024d); Wen et al. (2024b) also discuss certain topics related to the LLM knowledge boundary. However, Li et al. (2024d) lack a clear and formalized definition, and Wen et al. (2024b) merely focus on the abstention strategy for handling knowledge limitation. These limitations hinder a thorough and nuanced understanding of the LLM knowledge boundary.
@@
Building on our proposed definition, we systematically review related research. Our survey is organized around three key research questions. First, we address RQ1: Why study knowledge boundaries?, by detailing the LLMs’ undesirable behaviors that stem from their unawareness of knowledge boundaries (§ 3). Next, we explore RQ2: How can knowledge boundaries be identified?, highlighting uncertainty, calibration and probing techniques to distinguish different knowledge types (§ 4). Furthermore, we investigate RQ3: How can issues caused by knowledge boundaries be mitigated?, summarizing strategies to enhance the task performance and foster desired behaviors for each knowledge type (§ 5).
Finally, we discuss the open challenges and prospective directions for advancing the understanding of the LLM knowledge boundary. First, we advocate for more comprehensive benchmarks to assess knowledge boundaries across various types of knowledge limitations. Second, we emphasize the need for the generalization of knowledge boundary identification across domains and consider the potential utilization of LLM knowledge boundaries in future developments of LLMs. Lastly, we address unintended side effects of mitigation strategies, including over-refusal of knowledge within the boundaries and unnecessary costs incurred by implementing these strategies. The overview of this survey and related datasets are presented in Appendix A and B, respectively.
@@ 3 Undesired Behaviours
We first address RQ1: Why study knowledge boundaries? Due to the unawareness of knowledge boundary, LLMs exhibit various undesired behaviors that compromise the reliability and utility of their outputs, including factuality hallucinations, untruthful responses misled by the context, and truthful but undesired responses, posing challenges for the successful applications of LLMs.
@@ 3.1 Factuality Hallucinations
Factuality hallucinations Huang et al. (2023a) in LLMs occur when the model output diverges from real-world facts, which typically stem from the deficiency in the model’s domain-specific knowledge, outdated knowledge encoded within the model, and overconfidence in addressing unknowns.
@@
3.2 Untruthful Responses Misled by Context
Even though LLMs possess the required knowledge, they often produce untruthful responses when misled by context, which occurs in two forms: untruthful context, where the context includes false or misleading information, and irrelevant context, where extraneous details divert the model from generating precise responses.
Untruthful Context
Incorporating false information into the context significantly biases LLMs, severely impacting their performance Chen et al. (2024a); Pan et al. (2023). Using in-context learning (ICL) allows for editing factual knowledge in LLMs, which may lead to varied factual outputs Zheng et al. (2023a). When faced with untruthful views, LLMs often fail to stay true, being swayed by persuasive tactics despite initially correct responses Wang et al. (2023a); Xu et al. (2024b).
Irrelevant Context
@@ 3.3 Truthful but Undesired Responses
LLMs sometimes produce accurate yet improper responses when handling certain knowledge, leading to answers misaligned with user expectations.
@@
Controversial knowledge involves subjective questions with varied answers depending on individual perspectives Wang et al. (2024e); Amayuelas et al. (2024). These reveal biases in LLMs trained on skewed datasets, leading to partiality in responses. Such bias may cause unfair emphasis on certain viewpoints or stereotypical portrayals of demographics, exacerbating disparities Singh et al. (2024); Naous et al. (2024).
@@ 4.1 Uncertainty Estimation
Uncertainty estimation (UE) aims to quantify the uncertainty of a model regading its predictions for a given input. High uncertainty indicates that the model is unlikely to produce correct predictions to the input, thus the input-related knowledge lies outside of certain knowledge boundaries of the model. UE has been widely studied on NLP models Hu et al. (2023). In the era of LLMs, we highlight the following three groups of studies.
@@
The uncertainty of LLM can be decomposed into epistemic uncertainty and aleatoric uncertainty Hou et al. (2024). Epistemic uncertainty refers to the model-specific uncertainty, quantifying the lack of model knowledge, which is related to our definition of Parametric Knowledge Boundary. Aleatoric uncertainty refers to the data-level uncertainty, such as ambiguous prompts having multiple valid answers, referring to the gap between Outward Knowledge Boundary and Parametric Knowledge Boundary. Quantifying these types of uncertainty can help to identify different approaches for mitigating the knowledge limitations (Section 5). Solutions to quantify the two types of uncertainty can be roughly classified into data-side and model-side approaches, where one type of uncertainty can be obtained by subtracting the other type from the total uncertainty. The data-side quantification include input-side clarification and perturbation Hou et al. (2024); Ling et al. (2024); Gao et al. (2024), and output-side variation estimation Yadkori et al. (2024); Aichberger et al. (2024). The model-side quantification include model parameter and configuration perturbation Ling et al. (2024) and model internal states perturbation Ahdritz et al. (2024).
@@
Stemming from the traditional UE, the straightforward token probability-based UE computes the average token probability or the entropy of the LLM predictions as the uncertainty Manakul et al. (2023); Huang et al. (2023b). Detailed designs involve considering different granularities of the predictions beyond token-level, such as sentence-level Duan et al. (2023) and atomic fact-level Fadeeva et al. (2024), weighted by the relevance of different components Duan et al. (2023).
@@
The token probability-based UE are unsuitable for proprietary LLMs, and might be insufficient in quantifying the semantic uncertainty of LLM predictions. Therefore, the semantic-based UE is proposed, roughly categorized into consistency-based methods and verbalized methods. The consistency-based methods view the inconsistency among multiple sampled predictions of the input as the uncertainty. The approaches to measure the semantic consistency of the sampled outputs include the semantic distance calculated by smaller models Kuhn et al. (2023); Lin et al. (2024); Zhao et al. (2024c); Nikitin et al. (2024); Manakul et al. (2023), and the consistency in the LLM evaluation Chen and Mueller (2024); Manakul et al. (2023). The verbalized methods aim to enable LLMs to express their uncertainty directly as output tokens. Zhou et al. (2024) reveal that LLMs are reluctant to verbally express their uncertainty, possibly related to the lack of uncertainty expression in the training data. Lin et al. (2022a) and Chaudhry et al. (2024) adopt ICL and fine-tuning approaches to teach LLMs to generate uncertainty expressions.
@@ 4.2 Confidence Calibration
4.2.1 Prompt-based Calibration
Prompting LLMs to elicit confidence
This approach generally uses the prediction probability as a measure of the LLM confidence, estimated by the frequency of the prediction among multiple sampled predictions of the same input Si et al. (2023); Wang et al. (2023b), or by the probability of the prediction being evaluated as correct by LLMs Kadavath et al. (2022). Techniques to improve calibration include prompt ensemble Jiang et al. (2023a), hybrid approach Chen and Mueller (2024), fidelity evaluation Zhang et al. (2024d), and model ensemble Shrivastava et al. (2024); Feng et al. (2024).
Prompting LLMs to express confidence
This approach enables LLMs to directly generate the confidence as tokens in the prediction. Prompting RLHF-LLMs to express confidence can achieve better calibration than using token probability Tian et al. (2023), and prompting LLMs to generating explanations can further be leveraged to enhance calibration Zhao et al. (2024b); Li et al. (2024c). Combination with the former prompting approach can further improve performance Xiong et al. (2024b).
4.2.2 Fine-tuning for Calibration
The fine-tuning methods involve self-updating the LLM parameters and tuning additional models for calibration. The self-update involves instruction tuning for confidence expression Tao et al. (2024), and learning to adjust the output token probabilities Liu et al. (2024d); Xie et al. (2024). Additional models can be trained for adjusting the LLM output probability towards calibration Shen et al. (2024), or directly evaluating the correctness and estimating the confidence level of the LLM outputs Mielke et al. (2022); Stengel-Eskin et al. (2024).
@@ 5.1 Prompt-sensitive Known Knowledge
5.1.1 Prompt Optimization
Optimizing the prompt phrasing is essential for the LLM knowledge utilization and improved task performance. The prompt optimization can be categorized into two areas: instruction optimization and demonstration optimization.
For instruction optimization, training-free methods include search-based techniques like Monte Carlo search Zhou et al. (2023b); Li et al. (2023a); Yang et al. (2024c), tree search Wang et al. (2024d), and searching on edit operations Prasad et al. (2023), where the LLM is often involved as the prompt optimizer Yang et al. (2024a); Pryzant et al. (2023); Long et al. (2024). The training-based methods typically rely on reinforcement learning to train additional modules for prompt optimization Zhang et al. (2023a); Deng et al. (2022); Diao et al. (2023).
@@ 5.1.1 Prompt Optimization
For demonstration optimization, the diversity and similarity of the demonstrations are crucial factors for optimization Xu et al. (2024c). For example, the similar demonstrations are found by K-Nearest Neighbors Liu et al. (2022a) and BM25 Luo et al. (2023), and the diverse demonstrations are identified by support example selection Li and Qiu (2023) and diversity sampling Mavromatis et al. (2023). Effective demonstrations can also be identified by training ranking models according to better LLM task performance Li et al. (2023c); Rubin et al. (2022); Iter et al. (2023); Ye et al. (2023).
5.1.2 Prompt-based Reasoning
Prompt-based reasoning strategies are often adopted to improve the LLM knowledge utilization Wei et al. (2022); Zhou et al. (2023a); Yao et al. (2023); Zheng et al. (2023b). For multi-step knowledge-based QA, the process generally involves individual steps such as question decomposition Press et al. (2023), knowledge elicitation and inference Wang et al. (2022); Jung et al. (2022); Liu et al. (2022b). External knowledge is often involved in this process to mitigate the knowledge gaps Zhang et al. (2024c); Wu et al. (2024a); Zhao et al. (2023); Li et al. (2024e); Trivedi et al. (2023).
5.1.3 Self-refinement
The iterative self-refinement of the initial LLM prediction is also beneficial for knowledge utilization. The approaches can be broadly divided into single-model refinement and multi-agent debate. For single-model refinement, LLMs are prompted to refine the predictions under a designed evaluation and regeneration process Madaan et al. (2024); Miao et al. (2024), or generate self-verification questions to check for prediction consistency Manakul et al. (2023); Weng et al. (2023); Dhuliawala et al. (2024). While Huang et al. (2024) critique that LLMs struggle to achieve self-refinement without external feedback, Li et al. (2024b) show that self-estimated confidence may improve self-refinement. In multi-agent debate, the LLM plays different roles to assess and refine its predictions from multiple angles Du et al. (2024); Fu et al. (2023).
5.1.4 Factuality Decoding
Different decoding strategies can also affect the LLM knowledge utilization, thus affecting the prediction factuality, which falls into two categories Bi et al. (2024). The first category involves contrastive decoding against naive predictions with potential factual errors. The predictions for contrast come from smaller LLMs Li et al. (2023b), lower layers of the LLM Chuang et al. (2024); Chen et al. (2024b), tokens with lower predicted probabilities Kai et al. (2024), or predictions with induced hallucination Yang et al. (2024b); Zhang et al. (2023b). The second category leverages the truthful directions identified from LLM internal states (§ 4.3). By editing these internal representations during decoding, it steers the model towards truthful directions, thereby enhancing the factuality of predictions Li et al. (2024a); Chen et al. (2024e); Qiu et al. (2024); Chen et al. (2024g); Zhang et al. (2024e).

### v2

@@ Abstract
Although large language models (LLMs) store vast amount of knowledge in their parameters, they still have limitations in the memorization and utilization of certain knowledge, leading to undesired behaviors such as generating untruthful and inaccurate responses. This highlights the critical need to understand the knowledge boundary of LLMs, a concept that remains inadequately defined in existing research. In this survey, we propose a comprehensive definition of the LLM knowledge boundary and introduce a formalized taxonomy categorizing knowledge into four distinct types. Using this foundation, we systematically review the field through three key lenses: the motivation for studying LLM knowledge boundaries, methods for identifying these boundaries, and strategies for mitigating the challenges they present. Finally, we discuss open challenges and potential research directions in this area. We aim for this survey to offer the community a comprehensive overview, facilitate access to key issues, and inspire further advancements in LLM knowledge research. Equal contribution. Work was done during an internship at SMU. Corresponding author.
@@
Moxin Li1∗   Yong Zhao12∗†   Wenxuan Zhang3   Shuaiyi Li4   Wenya Xie5 See-Kiong Ng1   Tat-Seng Chua1   Yang Deng2‡ 1National University of Singapore  2Singapore Management University 3Singapore University of Technology and Design 4The Chinese University of Hong Kong  5University of Minnesota - Twin Cities {limoxin,yzhao}@u.nus.eduydeng@smu.edu.sg
@@ 1 Introduction
We identify the major limitations in existing definitions of the LLM knowledge boundary. Firstly, the Know-Unknow Quadrant Yin et al. (2023); Amayuelas et al. (2024); Li et al. (2025) categorizes knowledge based on the LLM’s possession and the LLM’s awareness of such knowledge, but this definition is conceptual and lacks formalization. Besides, Yin et al. (2024) introduce a formalized definition separating the influence of the prompt from the LLM’s mastery of the knowledge, yet they merely focus on the knowledge boundary of a specific LLM which lacks comprehensiveness. Additionally, some recent surveys Li et al. (2024e); Wen et al. (2024b) also discuss certain topics related to the LLM knowledge boundary. However, Li et al. (2024e) lack a clear and formalized definition, and Wen et al. (2024b) merely focus on the abstention strategy for handling knowledge limitation. These limitations hinder a thorough and nuanced understanding of the LLM knowledge boundary.
@@
Building on our proposed taxonomy, we systematically review related research. Our survey is organized around three key research questions. First, we address RQ1: Why study knowledge boundaries?, by detailing the LLMs’ undesirable behaviors that stem from their unawareness of knowledge boundaries (§ 3). Next, we explore RQ2: How can knowledge boundaries be identified?, highlighting uncertainty, calibration and probing techniques to distinguish different knowledge types (§ 4). Furthermore, we investigate RQ3: How can issues caused by knowledge boundaries be mitigated?, summarizing strategies to enhance the task performance and foster desired behaviors for each knowledge type (§ 5).
Finally, we discuss the open challenges and prospective directions for advancing the understanding of the LLM knowledge boundary (§ 6). First, we advocate for more comprehensive benchmarks to assess knowledge boundaries across various types of knowledge limitations. Second, we emphasize the potential utilization of LLM knowledge boundaries in future developments of LLMs. Lastly, we discuss the role of knowledge boundary in different knowledge mechanisms.
The overview of this survey [^1] https://github.com/li-moxin/knowledge_boundary_survey. and related datasets are presented in Appendix A and B, respectively.
@@ 3 Undesired Behaviours
We first address RQ1: Why study knowledge boundaries? Due to the unawareness of knowledge boundaries, LLMs exhibit various undesired behaviors that compromise the reliability and utility of their outputs, posing challenges for the successful applications of LLMs. In this section, we delineate three types of undesired behaviours according to the unawareness of each knowledge boundary.
3.1 Untruthful Responses Misled by Context
Even though LLMs possess the required knowledge, they often produce untruthful responses when misled by context, i.e., the prompt-sensitive known knowledge. This can occur in two forms: untruthful context, where the context includes false or misleading information, and irrelevant context, where extraneous details divert the model from generating precise responses.
Untruthful Context
Incorporating false information into the context significantly biases LLMs, severely impacting their performance Chen et al. (2024a); Pan et al. (2023). Using in-context learning (ICL) allows for editing factual knowledge in LLMs, which may lead to varied factual outputs Zheng et al. (2023a). When faced with untruthful views, LLMs often fail to stay true Zhao et al. (2025), being swayed by persuasive tactics despite initially correct responses Wang et al. (2023a); Xu et al. (2024b).
Irrelevant Context
@@ 3.2 Factuality Hallucinations
Factuality hallucinations Huang et al. (2023b), i.e., the model output diverges from real-world facts, typically stem from the following causes, relating to the model-specific unknown knowledge.
@@ 3.3 Truthful but Undesired Responses
LLMs may produce improper responses when handling model-agnostic unknown knowledge, leading to answers misaligned with user expectations.
@@
Controversial knowledge involves subjective questions with varied answers depending on individual perspectives Wang et al. (2024f); Amayuelas et al. (2024). These reveal biases in LLMs trained on skewed datasets, leading to partiality in responses. Such bias may cause unfair emphasis on certain viewpoints or stereotypical portrayals of demographics, exacerbating disparities Singh et al. (2024); Naous et al. (2024).
@@ 4.1 Uncertainty Estimation
Uncertainty estimation (UE) aims to quantify the uncertainty of a model regarding its predictions for a given input. High uncertainty indicates that the model is unlikely to produce correct predictions to the input, thus the input-related knowledge lies outside of certain knowledge boundaries of the model. UE has been widely studied on NLP models Hu et al. (2023). In the era of LLMs, we highlight the following four groups of studies.
@@
The uncertainty of LLM can be decomposed into epistemic uncertainty and aleatoric uncertainty Hou et al. (2024). Epistemic uncertainty refers to the model-specific uncertainty, quantifying the lack of model knowledge, which is related to our definition of Parametric Knowledge Boundary. Aleatoric uncertainty refers to the data-level uncertainty, such as ambiguous prompts having multiple valid answers, referring to the gap between Outward Knowledge Boundary and Parametric Knowledge Boundary. Quantifying these types of uncertainty can help to identify different approaches for mitigating the knowledge limitations (Section 5). Solutions to quantify the two types of uncertainty can be roughly classified into data-side and model-side approaches, where one type of uncertainty can be obtained by subtracting the other type from the total uncertainty. The data-side quantification include input-side clarification and perturbation Hou et al. (2024); Ling et al. (2024); Gao et al. (2024b), and output-side variation estimation Yadkori et al. (2024); Aichberger et al. (2024). The model-side quantification include model parameter and configuration perturbation Ling et al. (2024) and model internal states perturbation Ahdritz et al. (2024).
@@
Conformal Prediction
Conformal Prediction Law (2006) quantifies the uncertainty of model outputs by identifying a set of outputs with a guaranteed probability that the correct output is included within the set. This approach offers advantages such as being logit-free and suitable for black-box LLMs Su et al. (2024). Several studies have explored and attempted to address the issue of overconfidence in LLM conformal prediction Ravfogel et al. (2023); Ye et al. (2024). Furthermore, conformal prediction has been applied in techniques such as prompt selection Zollo et al. (2024), decoding stopping rules for guaranteed generation Quach et al. (2024), and ensuring reliability in retrieval-augmented generation Li et al. (2024d).
@@
Stemming from the traditional UE, the straightforward token probability-based UE computes the average token probability or the entropy of the LLM predictions as the uncertainty Manakul et al. (2023); Huang et al. (2023c). Detailed designs involve considering different granularities of the predictions beyond token-level, such as sentence-level Duan et al. (2023) and atomic fact-level Fadeeva et al. (2024), weighted by the relevance of different components Duan et al. (2023).
@@
The token probability-based UE are unsuitable for proprietary LLMs, and might be insufficient in quantifying the semantic uncertainty of LLM predictions. Therefore, the semantic-based UE is proposed, roughly categorized into consistency-based methods and verbalized methods. The consistency-based methods view the inconsistency among multiple sampled predictions of the input as the uncertainty. The approaches to measure the semantic consistency of the sampled outputs include the semantic distance calculated by smaller models Kuhn et al. (2023); Lin et al. (2024b); Zhao et al. (2024c); Nikitin et al. (2024); Manakul et al. (2023), and the consistency in the LLM evaluation Chen and Mueller (2024); Manakul et al. (2023). The verbalized methods aim to enable LLMs to express their uncertainty directly as output tokens. Zhou et al. (2024) reveal that LLMs are reluctant to verbally express their uncertainty, possibly related to the lack of uncertainty expression in the training data. Lin et al. (2022a) and Chaudhry et al. (2024) adopt ICL and fine-tuning approaches to teach LLMs to generate uncertainty expressions.
@@ 4.2 Confidence Calibration
Prompt-based Calibration
One group of approaches aims to prompt LLMs to elicit confidence, according to the prediction probability as a measure of the LLM confidence via sampling Si et al. (2023); Wang et al. (2023b), or by the probability of the prediction being evaluated as correct by LLMs Kadavath et al. (2022). Techniques to improve calibration include prompt ensemble Jiang et al. (2023a), hybrid approach Chen and Mueller (2024), fidelity evaluation Zhang et al. (2024d), and model ensemble Shrivastava et al. (2024); Feng et al. (2024).
Another group of approaches aims to prompt LLMs to directly express confidence as tokens in the prediction. Prompting RLHF-LLMs to express confidence can achieve better calibration than using token probability Tian et al. (2023), and prompting LLMs to generating explanations can further be leveraged to enhance calibration Zhao et al. (2024b); Li et al. (2024c). Combination with the former prompting approach can further improve performance Xiong et al. (2024b).
Fine-tuning for Calibration
The fine-tuning methods involve self-updating the LLM parameters and tuning additional models for calibration. The self-update involves instruction tuning for confidence expression Tao et al. (2024), and learning to adjust the output token probabilities Liu et al. (2024c); Xie et al. (2024). Additional models can be trained for adjusting the LLM output probability towards calibration Shen et al. (2024), or directly evaluating the correctness and estimating the confidence level of the LLM outputs Mielke et al. (2022); Stengel-Eskin et al. (2024).
@@ 5.1 Prompt-sensitive Known Knowledge
Prompt Optimization
Optimizing the prompt phrasing is essential for the LLM knowledge utilization and improved task performance. This approach can be categorized into two areas: instruction optimization and demonstration optimization.
For instruction optimization, training-free methods include search-based techniques like Monte Carlo search Zhou et al. (2023b); Li et al. (2023b); Yang et al. (2024c), tree search Wang et al. (2024e), and searching on edit operations Prasad et al. (2023), where the LLM is often involved as the prompt optimizer Yang et al. (2024a); Pryzant et al. (2023); Long et al. (2024). The training-based methods typically rely on reinforcement learning to train additional modules for prompt optimization Zhang et al. (2023a); Deng et al. (2022); Diao et al. (2023).
@@
For demonstration optimization, the diversity and similarity of the demonstrations are crucial factors for optimization Xu et al. (2024c). For example, the similar demonstrations are found by K-Nearest Neighbors Liu et al. (2022a) and BM25 Luo et al. (2023), and the diverse demonstrations are identified by support example selection Li and Qiu (2023) and diversity sampling Mavromatis et al. (2023). Effective demonstrations can also be identified by training ranking models according to better LLM task performance Li et al. (2023d); Rubin et al. (2022); Iter et al. (2023); Ye et al. (2023).
Prompt-based Reasoning
Prompt-based reasoning strategies are often adopted to improve the LLM knowledge utilization Wei et al. (2022b); Zhou et al. (2023a); Yao et al. (2023); Zheng et al. (2023b). For multi-step knowledge-based QA, the process generally involves individual steps such as question decomposition Press et al. (2023), knowledge elicitation and inference Wang et al. (2022); Jung et al. (2022); Liu et al. (2022b). External knowledge is often involved in this process to mitigate the knowledge gaps Zhang et al. (2024c); Wu et al. (2024a); Zhao et al. (2023); Li et al. (2024f).
Self-refinement
The iterative self-refinement of the initial LLM prediction is also beneficial for knowledge utilization. The approaches can be broadly divided into single-model refinement and multi-agent debate. For single-model refinement, LLMs are prompted to refine the predictions under a designed evaluation and regeneration process Madaan et al. (2024); Miao et al. (2024), or generate self-verification questions to check for prediction consistency Manakul et al. (2023); Weng et al. (2023). While Huang et al. (2024b) critique that LLMs struggle to achieve self-refinement without external feedback, Li et al. (2024b) show that self-estimated confidence may improve self-refinement. In multi-agent debate, the LLM plays different roles to assess and refine its predictions from multiple angles Du et al. (2024); Fu et al. (2023).
Factuality Decoding
Different decoding strategies can also affect the LLM knowledge utilization, thus affecting the prediction factuality, which falls into two categories Bi et al. (2024). The first category involves contrastive decoding against naive predictions with potential factual errors Kim et al. (2024b). The predictions for contrast come from smaller LLMs Li et al. (2023c), lower layers of the LLM Chuang et al. (2024); Chen et al. (2024b), tokens with lower predicted probabilities Kai et al. (2024), or predictions with induced hallucination Yang et al. (2024b); Zhang et al. (2023b). The second category leverages the truthful directions identified from LLM internal states (§ 4.3). By editing these internal representations during decoding, it steers the model towards truthful directions, thereby enhancing the factuality of predictions Li et al. (2024a); Chen et al. (2024e); Qiu et al. (2024); Chen et al. (2024g); Zhang et al. (2024e).